OpenStudy (anonymous):

Factor the trinomial below. x2 - 13x + 40

sam (.sam.):

(x-5)(x-8)

OpenStudy (anonymous):

(x-5)(x-8)

OpenStudy (muhammad_nauman_umair):

=x^2-8x-5x+40 =x(x-8)-5(x-8) =(x-5)(x-8)